OpenStudy (anonymous):

(x+3)/(x+1)-1/x=1 I tried to move forward with common denominators of (x)(x+1) but then I got lost

OpenStudy (anonymous):

because then I got x^2 +2x-2=0 ??

OpenStudy (amistre64):

\[\frac{x+3}{x+1} - \frac{1}{x} = 1 ??\]

OpenStudy (anonymous):

final answer is x+1/x

OpenStudy (amistre64):

x^2 +3x -x -1 = x^2 +x

OpenStudy (anonymous):

not right

OpenStudy (amistre64):

we get: x-1 = 0 x = 1

OpenStudy (anonymous):

I forgot my negative...ok let me focus on my problems, lol

OpenStudy (amistre64):

;)

OpenStudy (anonymous):

x^2 +3x -x -1 = x^2 +x ...you don't still need the denominator on the left here?